OpenStudy (anonymous):

Solve the system of elimation -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 Need All Steps Written Out

OpenStudy (unklerhaukus):

if you add the third equation to either of the first two , the x's will cancel out

OpenStudy (anonymous):

Okay what do i do after that

OpenStudy (unklerhaukus):

What did you get?

OpenStudy (anonymous):

2y+3z=0 and -y+z=-3 Right?

OpenStudy (unklerhaukus):

Add the third equation to the first equation like this [2x+3y+3z = 5 ] + (-2x+2y+3z = 0 ) [2x+3y+3z] + (-2x+2y+3z) = [5] + (0)

OpenStudy (anonymous):

nvm it would be 3y+3z=5 and -y=z+-3 Right

OpenStudy (anonymous):

oh okay then what do i do

OpenStudy (unklerhaukus):

[2x+3y+3z] + (-2x+2y+3z) = [5] + (0) 2x+3y+3z -2x+2y+3z = 5 + 0 simplify this, hopefully all the x's will go

OpenStudy (anonymous):

so after you take out the xs your left with 3y+3z+2y+5z=5+0 Right

OpenStudy (unklerhaukus):

simplify

OpenStudy (anonymous):

then your left with 1y+2z=5 Right?

OpenStudy (unklerhaukus):

try that again 2x+3y+3z -2x+2y+3z = 5 + 0 (2-2)x + (3+2)y + (3+3)z = 5 + 0

OpenStudy (anonymous):

0x+1y+6z=5

OpenStudy (unklerhaukus):

almost (2-2)x + (3+2)y + (3+3)z = 5 + 0 0x + (3+2)y + 6z = 5